### v1

# columns=colsep=4pt, rows=rowsep=.5pt

Abstract
Prototypical self-supervised learning methods consistently suffer from partial prototype collapse, where multiple prototypes converge to nearly identical representations. This undermines their central purpose—providing diverse and informative targets to guide encoders toward rich representations—and has led practitioners to over-parameterize prototype sets or add ad-hoc regularizers, which mitigate symptoms rather than address the root cause. We empirically trace the collapse to the joint optimization of encoders and prototypes, which encourages a type of shortcut learning: early in training prototypes drift toward redundant representations that minimize loss without necessarily enhancing representation diversity. To break the joint optimization, we introduce a fully decoupled training strategy that learns prototypes and encoders under separate objectives. Concretely, we model prototypes as a Gaussian mixture updated with an online EM-style procedure, independent of the encoder’s loss. This simple yet principled decoupling eliminates prototype collapse without explicit regularization and yields consistently diverse prototypes and stronger downstream performance.
1 Introduction
Prototypical self-supervised learning (SSL) (Siméoni et al., 2025; Caron et al., 2021) have come to rival the effectiveness of language-supervised alternatives in representation learning (Fan et al., 2025). Yet recent work has demonstrated that many prototypical SSL approaches suffer from a phenomenon known as partial prototype collapse, in which multiple prototypes converge to indistinguishable representations (Govindarajan et al., 2023, 2024). This phenomenon could explain why over-parameterizing the number of prototypes used has been a popular choice to improve performance in recent prototypical frameworks (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025).
Our experiments show that the practice of jointly optimizing the encoder and prototypes contributes to this collapse. The intuition is that joint optimization drives prototypes towards redundant representations early in training by exploiting shortcuts that minimize the loss at the cost of diverse and semantically meaningful representations. This leads to prototype over-parametrization; an expensive solution that mitigates symptoms without directly addressing the underlying issues (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025) while impeding a model’s ability to learn stronger representations and diminishing robustness to imbalanced data distributions (Govindarajan et al., 2024; Wen et al., 2024).
We find that these effects are particularly pronounced in the instance-level formulations of prototypical SSL. In Fig. LABEL:fig:unif-dino, we reaffirm previous findings (Govindarajan et al., 2024), showing how DINO’s (Caron et al., 2021) final prototype distribution tends to collapse to one or two modes, which severely limits the diversity of representations. This phenomenon also persists in more recent methods such as DINOv2 (Oquab et al., 2024), cf. Fig. LABEL:fig:unif-dinov2. While recent efforts have looked to address this issue, either via explicit regularization (Govindarajan et al., 2024), cf. Fig. LABEL:fig:unif-dino-kp, or implicit mechanisms (Silva and Ramírez Rivera, 2023), cf. Fig. LABEL:fig:unif-carp, these models still exhibit collapse, suggesting that the underlying issue is not yet fully resolved.
Motivated by these observations, we introduce a decoupling strategy designed to break the joint-optimization procedure. In our approach, the encoder and prototypes are learned under separate objectives—see Fig. 2. Concretely, we model the prototypes as a Gaussian mixture, updated via an expectation maximization (EM) independently of the encoder’s loss. This decoupled training eliminates partial prototype collapse, cf. Fig. LABEL:fig:unif-carp-dec, and points to joint optimization as the most likely culprit driving partial prototype collapse.
Our main contributions are as follows: (i) we conduct a systematic and quantitative analysis of a broad range of prototypical SSL frameworks, demonstrating that partial prototype collapse extends well beyond the DINO family of models; (ii) we identify the underlying mechanism driving this collapse as the joint optimization of encoders and prototypes under a shared loss, which encourages redundant prototype representations; and (iii) we introduce a fully decoupled training framework that isolates prototype estimation from encoder learning, achieving consistently high prototype diversity throughout training and yielding stronger representations with improved robustness to long-tailed data distributions.
Figure 1: Uniformity of prototype representations. Prototypes are projected to $\mathbb{R}^{2}$ using principal component analysis (PCA). The resulting prototype distributions are visualized using (top) Gaussian kernel density estimation (KDE), and (bottom) the angular distributions are estimated with a von Mises-Fisher KDE with $\kappa=20$.
2 Understanding Partial Prototype Collapse
2.1 Preliminaries
Prototypical Self-Supervised Formulation. Among recent SSL approaches, prototypical formulations (where network outputs are aligned to a set of learnable, class-agnostic prototypes) have achieved state-of-the-art performance on several vision benchmarks (Siméoni et al., 2025). We begin by describing the general prototypical framework used throughout this paper. Our exposition emphasizes the instance-level objective (Caron et al., 2021; Silva and Ramírez Rivera, 2022, 2023), but the same formulation can readily be extended to incorporate a dense objective (Zhou et al., 2022) or reformulated with an explicit dense prediction objective (Darcet et al., 2025).
Given $J$ stochastic augmentations $\{v^{j}\}_{j=1}^{J}$ of an image $x$, a student backbone with projection head $f_{\theta}$ and an EMA-updated teacher $f_{\phi}$ map each view to a latent representation $h^{j}_{(\cdot)}=f_{(\cdot)}(v^{j})\in\mathbb{R}^{D}$. Similarity scores with a learnable prototype set $C=[c_{1},c_{2},\dots,c_{K}]\in\mathbb{R}^{D\times K}$ are computed as $z_{(\cdot)}^{j}=h_{(\cdot)}^{j}C^{\top}\in\mathbb{R}^{K}$ and converted into prototype-assignment probabilities via a softmax with an optional temperature parameter $\tau>0$ through
$p\left(z_{(\cdot)}^{j}\right)=\mathrm{softmax}\!\left(\frac{z_{(\cdot)}^{j}}{\tau}\right).$ (1)
The overall objective of prototypical frameworks is to enforce cross-view consistency between the student and teacher branches using a consistency loss $\mathcal{L}_{f}$ [^1] Common choices include the cross-entropy loss (Caron et al., 2021) and the cluster loss (Silva and Ramírez Rivera, 2023).—see Fig. LABEL:fig:trad-jea.
Partial Prototype Collapse. Govindarajan et al. (2024) investigated the DINO-family of methods (Caron et al., 2021; Zhou et al., 2022; Assran et al., 2022, 2023a) and demonstrated that these prototypical SSL approaches are susceptible to partial prototype collapse. This phenomenon arises when multiple prototypes converge to nearly identical representations during training. While this behavior does not amount to a complete collapse of all prototypes, it, nevertheless, reduces the diversity of representations and may hinder the effectiveness of downstream tasks. To formalize this behavior, the authors introduced the following definition of partial collapse.
Definition 2.1 (Partial prototype collapse).
Consider the set $C=\{c_{k}:k=1,\dots,K\}$ of $K$ prototype vectors, $c_{k}$ such that $\|c_{k}\|=1$. A partial prototype collapse (of degree $M$ and $\epsilon$ distance) is said to have occurred if there exists a set of $M$ disjoint partitions of prototype vectors $V_{m}\subset C$, $m=1,\dots,M$, and $M$ representative prototype vectors $v_{m}\in V_{m}$, such that for all $m=1,\dots,M$, $1-v_{m}^{\mathsf{T}}c_{j}<\epsilon$, for all $c_{j}\in V_{m}$. The set of $M$ unique prototypes is defined as $U=\{v_{m}\}_{m=1}^{M}$.
Applying this definition, they further showed that the DINO-family of methods retained only about 2%–40% of their initially initialized prototypes as unique, thereby, revealing a substantial redundancy in the learned representations. To alleviate this issue, they proposed applying KoLeo regularization (Beirlant et al., 1997; Delattre and Fournier, 2017; Sablayrolles et al., 2019) directly to the prototypes, effectively introducing a diversity-enforcing auxiliary term into the objective. While this helped reduce collapse, it did not fully solve the problem. Their approach also introduces a new hyperparameter, creating a delicate trade-off: insufficient regularization limits diversity, whereas excessive regularization impair encoder learning. While the work of Govindarajan et al. (2024) represented an important first step in identifying partial prototype collapse, their analysis did not examine its underlying mechanisms and was limited to methods within the DINO family. This opens the door to a broader line of inquiry: Is partial prototype collapse restricted to the DINO family of methods, and what mechanisms drive this type of collapse?
2.2 Do all prototypical SSL formulations exhibit partial prototype collapse?
Table 1: We evaluate the number of unique prototypes according to Definition 2.1 with $\epsilon=0.025$. Vanilla iBOT uses the same head for both objectives; for clarity, only one is shown.
Model Objective Init. Protos. Unique Protos. Dense Unique Protos. Instance (% of Init. protos.)
DINO Instance 60000 908 (1.5%)
CARP Instance 65536 7052 (10.8%)
CAPI Dense 16384 16383 (99.9%)
iBOT Hybrid 8192 $3057^{\star}$ (37.3%)
iBOT-vMF + KP Hybrid 8192 7895 (96.4%)
DINOv2 Hybrid 262144 110201 2556 (43.0% )
Despite recent progress, our understanding of prototype collapse remains incomplete. The evidence so far comes exclusively from the DINO family, leaving open whether its partial collapse reflects a universal tendency of prototypical SSL or an artifact of its particular architecture and training regime. To bridge this gap, we next investigate alternative prototypical frameworks, evaluating whether their learned prototypes exhibit similar degrees of redundancy—or if some, by design, avoid collapse altogether.
[…]
To our surprise, we find that CAPI (Darcet et al., 2025) maintains markedly higher prototype diversity than the other methods, exhibiting near-complete uniqueness with only one collapsed prototype out of 16,384. It even manages to surpass iBOT combined with the KoLeo-Proto (KP) regularization (Govindarajan et al., 2024), which explicitly encourages prototype diversity. This begs the question, what underlying mechanism in CAPI allows the prototypes to remain diverse throughout training?
A key distinction of CAPI compared to earlier prototypical joint-embedding frameworks is that it partially decouples the prototypes from the main loss. In most prototypical formulations—particularly within the DINO family—the prototypes used to compute the teacher’s soft assignments are an EMA-updated version of those used for the student, keeping teacher and student prototypes tightly coupled through a joint optimization objective. However, Darcet et al. (2025) argue that, in a masked image modeling (MIM) setting such as iBOT (Zhou et al., 2022), this coupling creates a distributional mismatch: the teacher’s prototype assignments are computed on unmasked patches, whereas the student’s prototype assignments are computed on masked patches. They argue that this joint optimization under distributional mismatch induces training instabilities.
CAPI addresses the instabilities driven by distributional mismatch, by introducing a partial decoupling strategy: the teacher encoder produces latent patch embeddings that are assigned to prototypes learned independently within a separate clustering module, and these assignments then serve as targets for the student. The student branch, in turn, predicts these assignments using its own prototypes, which continue to be updated jointly with the encoder. Although Darcet et al. (2025) proposed this mechanism primarily for stabilizing purposes, our analysis shows that it also alleviates partial prototype collapse to an extent and, as a result, improves prototype spread.
3 Solving Partial Prototype Collapse through Decoupling
Figure 2: (a) Traditional joint-embedding architectures with a prototypical formulation: encoders and prototypes are optimized jointly under the same loss, which can lead to shortcut learning and prototype collapse—prototypes converge to similar representations, reducing the effective representation space. (b) Our proposed solution: decouples the gradient flow to the prototypes and updates them with a separate objective, mitigating shortcut learning and preserving prototype diversity.
Our analysis of CAPI reveals that decoupling the teacher’s prototype updates from the main objective coincides with improved prototype diversity. This finding leads us to hypothesize that partial prototype collapse in prototypical SSL arises primarily from updating prototypes and encoders under a shared loss. When prototypes are optimized jointly with the encoder, they may drift toward similar representations that minimize prediction error without improving the underlying features—a form of shortcut learning. [^4] We clarify that by “mitigating shortcut learning” we do not mean shortcuts in relation to spurious correlations and group robustness (Geirhos et al., 2020), but to the early drift of prototypes into redundant representations that artificially reduce the loss without improving the encoder’s representations. Conversely, decoupling the teacher’s prototypes from the main loss, as in CAPI (Darcet et al., 2025), should reduce this incentive, leading to more stable and diverse prototype usage throughout training. This observation leads to our formal problem statement:
Problem Formulation. Traditional prototypical SSL methods jointly optimize an encoder $f_{\theta}$ and a set of prototypes $C=\{c_{k}\}_{k=1}^{K}$ by minimizing a consistency loss over augmented views,
$\min_{\theta,\,C}\;\mathcal{L}_{f}\bigl(f_{\theta},\,C\bigr).$ (2)
This joint optimization often induces a form of shortcut learning, where the prototypes’ distribution rapidly collapse into a narrow region of the representation space early in training. Such premature collapse undermines the very purpose of learning the prototypes CC—to provide diverse and informative targets that guide $f_{\theta}$ toward representations that transfer well to downstream tasks.
If partial prototype collapse truly stems from the joint optimization of prototypes and encoders, then separating their updates should prevent the collapse. Motivated by this intuition, and unlike CAPI (Darcet et al., 2025), which decouples only the teacher’s prototypes from the main loss while keeping the student prototypes jointly optimized, we propose a fully decoupled training procedure that isolates prototype estimation from encoder learning, cf. Fig. LABEL:fig:dec-jea, allowing us to directly probe the role of joint optimization in driving collapse.
Proposed Solution: Full Decoupling. Instead of jointly solving the original loss (2), over the iterations $t$, we alternate two separate objectives: (i) update the prototypes by solving an independent unsupervised estimation problem on the latent features, i.e.,
$C^{t}=\operatorname*{arg\,min}_{C\in\mathcal{C}}\;\mathcal{L}_{C}\left(C^{t-1};h^{t}_{\phi}\right),$ (3)
and (ii) update the encoder for fixed prototypes by minimizing the consistency loss, i.e.,
$\theta^{t+1}=\operatorname*{arg\,min}_{\theta}\;\mathcal{L}_{f}\left(h^{t},C^{t}\right),$ (4)
where $h^{t}$ denotes the latent representations at iteration $t$, and $\mathcal{L}_{C}$ is a loss that estimates prototypes directly from the latent features, independent of the encoder’s loss $\mathcal{L}_{f}$. By fully separating prototype estimation from encoder optimization, our method removes the shortcut incentive entirely and aims to achieve stable, diverse prototypes throughout training.
3.1 Decoupling the Optimization
A central question in our framework is how to estimate prototypes once they are decoupled from the encoder. Breaking the joint optimization opens up a rich design space: many unsupervised objectives could, in principle, be used for prototype estimation in Eq. (3). To be effective, however, such objectives must satisfy three key properties: (i) they should be representative and distinctive, ensuring that each prototype captures a coherent and separate mode of the data; (ii) they should be learned over the evolving dataset rather than isolated mini-batches, to avoid noisy estimates and the resulting training instabilities; and (iii) they should be computationally efficient, so as not to impede training time or create memory bottlenecks.
An obvious way to optimize Eq. (3) is to run $K$-Means clustering on the latent features. However, naively applying $K$-Means at every iteration violates two of the key properties outlined above: it bases prototype updates on the current mini-batch rather than the evolving dataset—breaking property (ii)—and it is prohibitively expensive to run online at scale—breaking property (iii). DeepCluster (Caron et al., 2018) and PCL (Li et al., 2021) alleviated the cost by clustering only once per epoch on the full dataset, but this left prototypes outdated relative to the changing representation space. SWaV (Caron et al., 2020) addressed the issue of outdated prototypes by introducing an online, gradient-based method for prototype updates; however, this introduced the very joint-optimization dilemma we seek to avoid.
We address the optimization problem of Eq. (3), while satisfying the three outlined properties, by representing the prototypes as the means of an online Gaussian Mixture Model (GMM) (Neal and Hinton, 1998; Sato and Ishii, 2000). In contrast to naive $K$-Means, the online GMM incrementally updates its mixture components as new data arrive, allowing the prototypes to evolve continuously with the representation space. This procedure naturally incorporates information from the entire dataset over time rather than relying on isolated mini-batches, thereby satisfying property (ii), while its incremental updates render it computationally feasible at scale, satisfying property (iii). Moreover, mixture-based clustering has already been successfully applied in several deep learning contexts (Pu et al., 2023; Liang et al., 2022; Zhao et al., 2023), underscoring its potential to satisfy property (i).
While incorporating the GMM into our framework, we developed a variant specifically designed to cope with the high-dimensional feature spaces and large prototype counts that naturally emerge in large-scale representation learning. Without modification, standard online updates at this scale suffer from unbalanced responsibilities and component drift. Drawing on responsibility-weighted forgetting (Celaya and Agostini, 2015) and deterministic annealing (Ueda and Nakano, 1998), we modulate both the strength of parameter updates and the sharpness of assignments. This ensures that components with fewer assignments maintain stable estimates while those with higher usage remain responsive, resulting in a more stable mixture and higher-quality prototypes at scale. We refer the reader to Appendix A for further implementation details.
4 Experiments
4.1 Experimental Setup
To evaluate our decoupling strategy, we adopt CARP (Silva and Ramírez Rivera, 2023). We select CARP for two main reasons. First, it is an instance-based approach, which we have shown to exhibit a higher degree of prototypical collapse (cf. Table 1) which offers a more challenging setting to evaluate the efficacy of our decoupling method, as the collapse is more pronounced, while also isolating the effect of MIM objectives on prototype diversity, which we leave for future work. Second, extensive ablation experiments have demonstrated that CARP maintains robust and stable performance across a wide range of hyperparameters (Silva and Ramírez Rivera, 2023), making it a suitable framework for evaluating our decoupling strategy without extensive hyperparameter tuning. [^5] This contrasts with the DINO formulation, which has been shown in several studies to suffer from substantial hyperparameter sensitivity (Ruan et al., 2023; Wu et al., 2025).
Except for the experiment reported in Section 4.2, where we rely on the officially released weights, we train all models ourselves. [^6] Details of how the baselines were trained are provided in Appendix B. This enables us to analyze the training dynamics of the different methods in greater depth (an analysis that would not be possible without access to intermediate checkpoints) and to provide a more detailed assessment of how prototype diversity benefits the handling of long-tailed distributions.
4.2 Uniqueness of prototypes under different thresholds
[…]
By examining Fig. LABEL:fig:unique-protos-v-eps, several notable insights emerge. First, a substantial collapse is observed in DINO and DINOv2, highlighting the extent of collapse within the DINO family of methods when no preventive measures are applied. Second, CAPI preserves prototype diversity at lower thresholds of $\epsilon$, exhibiting only minor collapse. As the constraint increases, CAPI still retains approximately $38\%$ of its initialized prototypes. We attribute CAPI’s robustness to prototypical collapse to its partial decoupling mechanism, in which the teacher-branch prototypes are updated using a loss function separate from that of the student-branch encoder.
Lastly, we observe no evidence of partial prototype collapse across all tested values of $\epsilon$ when the prototype objective is fully decoupled from the overall loss. This stands in clear contrast to CARP, which exhibits collapse in $90\%$ of its prototypes already at $\epsilon=0.025$ and CAPI with its partial decoupling. Crucially, our approach not only prevents partial prototype collapse but also improves the encoder’s learned representations, as demonstrated in Fig. LABEL:fig:carp-decoupled. This distinction is central to our contribution: it supports our claim that joint optimization introduces a form of shortcut learning (minimizing the loss without genuinely enriching the encoder’s representations) whereas decoupling avoids this pitfall and yields more informative features.
Figure 3: (a) Unique prototypes versus $\epsilon$ (Definition 2.1). At $\epsilon=0$ all initialized prototypes are counted as unique, whereas increasing $\epsilon$ enforces stricter criteria (e.g., at $\epsilon=0.5$ only prototypes separated by at least $60^{\circ}$ are considered unique). (b) k-NN performance on IN1k for CARP showcasing an improvement with decoupling.
4.3 Training dynamics
Although previous studies have examined partial prototype collapse, they have focused almost exclusively on its manifestation at the end of training (Govindarajan et al., 2023, 2024), leaving its interaction with the training dynamics largely unexplored. To address this and gain deeper insight into the relationship between prototype collapse and training progression, in Fig. 4, we track the number of unique prototypes in use and the linear evaluation accuracy on ImageNet-1k (Deng et al., 2009) over the first 100 epochs for multiple methods. Specifically, we include CARL (Silva and Ramírez Rivera, 2022) and CARP (Silva and Ramírez Rivera, 2023) (CARP being an extension of the CARL framework that introduces random partitioning) alongside DINO (Caron et al., 2021), DINO + KoLeoProto (Govindarajan et al., 2024), and our proposed decoupled approach.
A first observation is that partial prototype collapse emerges very early in training: after only 10 epochs, two-thirds of the prototypes have already collapsed. Introducing the KoLeo-Proto (KP) regularizer (Govindarajan et al., 2024) proves effective at maintaining prototype diversity, underscoring its role as a strong collapse-prevention mechanism. While DINO+KP exhibits a modest drop in accuracy during the early stages of training, this effect diminishes over time, and the final model slightly outperforms vanilla DINO. This suggests that preserving higher prototype diversity throughout training ultimately leads to stronger representations in the converged model.
Another noteworthy observation is that CARP, is far less prone to prototype collapse than the baseline CARL. By the end of training, CARL retains only about $0.2\%$ of its initialized prototypes as unique, whereas CARP preserves $9\%$. This difference is also reflected in final linear evaluation performance: CARP outperforms CARL late in training but is impaired early in training. While these gains cannot be attributed solely to prototype uniqueness (training stability likely plays a significant role) they nevertheless amount to a substantial margin of roughly $4$ percentage points in linear accuracy by the end of training.
Finally, we find that CARP + Decoupling maintains full prototype utilization while outperforming all other methods in the early and late phases of training, indicating that decoupling does not hinder early training dynamics. By the end of training, CARP + Decoupling improves upon CARP, demonstrating that enhanced prototype diversity yields stronger representations.
Figure 4: (a) Number of unique prototypes over epochs. Prototypical methods without explicit diversity mechanisms (e.g., DINO and CARL) exhibit substantial prototype collapse early in training. (b) Linear evaluation accuracy on ImageNet-1k. Higher prototype diversity generally correlates with stronger final performance, though some methods (e.g., DINO + KP) imposes a cost of slower early convergence.
4.4 Effect of prototype diversity across head, medium and tail classes
Prototypical SSL methods generally perform well on carefully curated datasets; however, their effectiveness deteriorates markedly when trained on uncurated data (Oquab et al., 2024; Siméoni et al., 2025), which commonly exhibit class imbalances and long-tailed distributions characteristic of real-world scenarios (Newman, 2005; Mahajan et al., 2018; Van Horn et al., 2018). We argue, therefore, that a desirable property of any pre-training framework is robustness to long-tailed data distributions. Such robustness would enable practitioners to improve learned representations simply by scaling the amount of training data. Nevertheless, Fan et al. (2025) report that scaling up the amount of training data for a 7B-parameter version of DINOv2 yielded limited benefits: as the size of the training data increased, performance across a range of benchmarks decreased. These findings on data scaling underscore a fundamental limitation of current prototypical SSL approaches.
Table 2: Performance comparison on the iNaturalist 2018 dataset. Results are reported as Top-1 accuracy (%) for different class splits.
Methods
Head
(>100>100)
Medium
(>20&100>20\ $\leq$ 100)
Tail
($\leq 20$)
All
DINO 55.2 46.2 41.9 45.3
DINO + KP 59.8 $\uparrow$  4.6 50.4 $\uparrow$  4.2 45.0 $\uparrow$  3.1 49.0 $\uparrow$  3,7
CARP 56.0 46.9 42.5 45.9
CARP + Decoupling 59.1 $\uparrow$  3.1 49.3 $\uparrow$  2.4 45.9 $\uparrow$  3.4 48.9 $\uparrow$  3.0
[…]
Table 3: ImageNet evaluation for instance-based prototypical SSL methods.
Method Backbone Epochs k-NN (%) Linear (%)
DINO RN-50 $400$ 67.5 75.3
SWAV RN-50 $400$ 65.0 74.6
DeepCluster-v2 RN-50 $800$ 66.6 75.2
CARP RN-50 $400$ 67.7 75.3
CARP + Decoupling RN-50 $400$ 69.1 75.3
DINO ViT-S/16 $300$ 72.8 76.2
CARP + Decoupling ViT-S/16 $300$ 74.1 76.2
MSN ViT-S/16 $600$ 76.9
DINO ViT-S/16 $800$ 74.5 77.0
DINO-vMF ViT-S/16 $800$ 74.7 77.0
CARP + Decoupling ViT-S/16 $800$ 75.3 76.4
DINO ViT-B/16 $400$ 76.1 78.2
DINO-vMF ViT-B/16 $400$ 77.4 78.8
CARP + Decoupling ViT-B/16 $400$ 76.7 78.1
Results computed by us using the officially released pre-trained models.
Official pre-trained weights are not publicly available
Table 2 shows that greater prototype diversity substantially enhances robustness in long-tailed datasets. While DINO exhibits a 3.7-percentage-point improvement attributable to the increase in prototype diversity achieved through KP regularization, our proposed decoupling approach on top of CARP yields an improvement of $3.0$ percentage points, although CARP already enjoys greater robustness on long-tailed data due to its higher prototype diversity than DINO. Furthermore, we find that all examined distributions of class frequencies benefit from increased prototype diversity. This finding underscores the necessity of mitigating prototypical collapse in prototypical SSL frameworks and suggests that integrating such mechanisms may enhance the robustness of methods trained on less curated data. Finally, CARP + Decoupling surpasses DINO + KP in tail-class accuracy by nearly one percentage point, this combined with the substantial gains obtained through decoupling the joint optimization in CARP, highlights the efficacy of our decoupling strategy in enhancing prototype diversity.
4.5 Linear Classification on ImageNet-1k
We obtain strong k-NN performance with our proposed decoupling approach built on top of CARP, demonstrating the benefit of increased prototype spread with a $1.8$-percentage-point improvement in k-NN performance (Table 3). As is well known (Caron et al., 2021), linear evaluation results are highly sensitive to hyperparameter choices, particularly for methods with diverse prototypes (Darcet et al., 2025; Oquab et al., 2024). Consequently, rather than performing extensive hyperparameter tuning for linear classification, we follow a light grid-search protocol [^7] See Appendix C.1 for an extended discussion of the linear evaluation protocol and additional results. and place primary emphasis on the k-NN metric, which provides a more robust, fine-tuning-free assessment of representation quality.
5 Related Work
Self-Supervised Learning leverages unlabeled data by enabling the model to generate its own supervisory signals. Early approaches introduced pretext tasks such as image inpainting (Pathak et al., 2016) and solving jigsaw puzzles (Noroozi and Favaro, 2016). Since then, a dominant paradigm has emerged that employs joint-embedding architectures within a contrastive framework (Chen et al., 2020, 2021; Chen and He, 2021; Grill et al., 2020), which was later extended by incorporating a predictor network (JEPA) (Assran et al., 2023b). Another line of work departs from joint-embedding architectures, using reconstruction-based objectives such as MAE (He et al., 2022) and BEiT (Bao et al., 2022) based on MIM objectives.
Prototypical Self-Supervised Learning extends joint-embedding methods by introducing a discrete set of learnable prototypes that serve as targets for representation assignment. Such formulations have achieved state-of-the-art performance among computer-vision–only SSL approaches (Siméoni et al., 2025; Venkataramanan et al., 2025), rivaling the effectiveness of language-supervised alternatives (Fan et al., 2025; Radford et al., 2021). Early methods performed cluster assignments once per epoch (Caron et al., 2018; Li et al., 2021), which could result in outdated prototypes. To mitigate this issue, online methods were introduced (Asano et al., 2019; Caron et al., 2020), enabling the joint optimization of the encoder and prototypes, a strategy that has since become the status quo for most prototypical formulations (Assran et al., 2022; Oquab et al., 2024; Siméoni et al., 2025; Silva and Ramírez Rivera, 2022; Caron et al., 2021; Zhou et al., 2022; Assran et al., 2023a; Ruan et al., 2023). However, this joint optimization also introduced the phenomenon of partial prototype collapse which Govindarajan et al. (2023) was first to identify, and has since been addressed to varying degrees of success both explicitly (Govindarajan et al., 2024; Wen et al., 2024) and implicitly (Silva and Ramírez Rivera, 2023; Darcet et al., 2025). Recent work departs from learning prototypes explicitly, instead using fixed high-dimensional codes sampled from a Rademacher distribution (Sansone et al., 2025) or non-parametric targets built from a queue of past encoder representations that act as effective prototypes (Gidaris et al., 2021, 2024; Silva et al., 2024, 2025).
6 Conclusion
This work provides, to the best of our knowledge, the first systematic investigation into the root causes of partial prototype collapse in prototypical self-supervised learning. Through an extensive empirical analysis across a diverse set of frameworks, we show that prototype collapse is not confined to the DINO family but is a widespread phenomenon, drastically reducing the effective diversity of prototypes and limiting downstream transfer performance. We identify joint optimization of encoders and prototypes under a shared loss as the key mechanism driving this collapse-encouraging a type of shortcut learning and redundant prototype representations early in training.
Building on this diagnosis, we introduce a fully decoupled training framework that isolates prototype estimation from encoder learning. By representing prototypes as a continuously updated Gaussian mixture and estimating them via an online EM-style procedure, our approach removes the shortcut incentive inherent to joint optimization. This design yields stable and highly diverse prototypes throughout training without ad-hoc regularization or hyperparameter trade-offs.
Across extensive experiments, including imbalanced datasets, our decoupling strategy consistently enhances prototype spread, strengthens learned representations, and improves robustness to challenging data distributions. Together, these findings highlight the importance of breaking the joint optimization paradigm and position decoupled prototype learning as a simple, principled, and scalable path forward for prototypical SSL.
Appendix A Online Gaussian Mixture Model
To decouple the joint optimization of the encoder and prototypes, we choose to represent prototypes explicitly as components of a Gaussian Mixture Model (GMM). Under this probabilistic formulation, prototypes capture the underlying structure of the data distribution independently from the encoder’s training objective. By leveraging an Expectation-Maximization (EM)-based approach (Dempster et al., 1977), prototypes maximize the data likelihood of the teacher’s latent embeddings, thereby, effectively disentangling their update mechanism from the encoder’s loss-driven gradient updates. This separation ensures that prototype optimization is solely governed by the statistical characteristics of the latent representations ($\mathcal{L}_{C}$), promoting robust learning and preventing partial collapse.
While the classical EM algorithm provides a principled framework for prototype learning, it assumes access to a static and complete dataset. This assumption is incompatible with most SSL frameworks, where data becomes available incrementally as the encoder continuously updates its representations during training. To address this limitation, we propose to adopt an online variant of the EM algorithm (Neal and Hinton, 1998; Sato and Ishii, 2000), which updates the mixture parameters incrementally using mini-batches of data.
At each iteration $t$, we treat the incoming batch as a sample from the evolving feature distribution and perform a two-step update: computing responsibilities for each latent vector, followed by an update of the GMM parameters using sufficient statistics.
The GMM is parameterized by $\Psi^{(t-1)}=\{\pi^{(t-1)},\mu^{(t-1)},\Sigma^{(t-1)}\}$, where $\pi$, $\mu$, and $\Sigma$ denote the mixture weights, means (prototypes), and diagonal covariances, respectively. For each input $i$, the responsibility $\gamma_{ik}^{(t)}$ is defined as:
$\gamma_{ik}^{(t)}=\frac{\pi_{k}^{(t-1)}P\left(h^{(t)}_{i}\>\lvert\>\mu_{k}^{(t-1)},\Sigma_{k}^{(t-1)}\right)^{\beta}}{\sum_{k^{\prime}=1}^{K}\pi_{k^{\prime}}^{(t-1)}P\left(h^{(t)}_{i}\>\lvert\>\mu_{k^{\prime}}^{(t-1)},\Sigma_{k^{\prime}}^{(t-1)}\right)^{\beta}}.$ (A.1)
Here, $\gamma_{ik}^{(t)}$ denotes the soft assignment of latent vector $h^{(t)}_{i}=f^{(t)}_{\phi}(v_{i})$ from the teacher network to the $k$-th Gaussian component. Intuitively, it reflects how well the $k$-th component’s distribution explains $h^{(t)}_{i}$. To further enhance prototype utilization, particularly when employing a large number of components $K$, we introduce an annealing mechanism controlled by a smoothing factor $\beta\in[0,1]$ (Ueda and Nakano, 1998), which modulates the sharpness of the responsibilities during training.
Following the assignment of responsibilities, the expectation step utilizes these values to compute the intermediate sufficient statistics for the current timestep necessary to update the mixture’s sufficient statistics. Specifically, at each time step, the intermediate sufficient statistics for a batch $B$ and views $J$ are calculated as:
$\tilde{S}^{(t)}_{k}=\left\{\tilde{S}_{k,\pi}^{(t)},\;\tilde{S}_{k,\mu}^{(t)},\tilde{S}_{k,\Sigma}^{(t)}\right\}=\left\{\sum_{i=1}^{JB}\gamma_{ik}^{(t)},\sum_{i=1}^{JB}\gamma_{ik}^{(t)}h_{i}^{(t)},\sum_{i=1}^{JB}\gamma_{ik}^{(t)}h_{i}^{(t)}h_{i}^{(t)^{T}}\right\}.$ (A.2)
These intermediate statistics are then used to update the sufficient statistics of the mixture using a weighted moving average:
$S_{k}^{(t)}=\begin{cases}\eta^{\hat{\gamma}_{k}}\,S_{k}^{(t-1)}+\bigl(1-\eta^{\hat{\gamma}_{k}}\bigr)\,\tilde{S}_{k}^{(t)},&\text{if }t>1,\\[2.0pt] \tilde{S}_{k}^{(1)},&\text{otherwise.}\end{cases}$ (A.3)
$\text{where }\quad\tilde{S}_{k}^{(1)}=\left\{\frac{JB}{K},\;\mu_{k}^{(1)}\,\tilde{S}_{k,\pi}^{(1)},\;\Sigma_{k}^{(1)}\,\tilde{S}_{k,\pi}^{(1)}+\frac{\tilde{S}_{k,\mu}^{(1)}\tilde{S}_{k,\mu}^{(1)\top}}{\tilde{S}_{k,\pi}^{(1)}}\right\}.$ (A.4)
Here, $\eta\in[0,1]$ is a forgetting factor, controlling the influence between new and past information. When using a large number of prototypes $K$ the mixture’s components will receive sparse updates, thus, the forgetting factor $\eta$ will tend to push the sufficient statistics towards zero. To circumvent this issue, we introduce a responsibility based forgetting mechanism (Celaya and Agostini, 2015) through the incorporation of the expectation of the responsibilities over the batch dimension $\hat{\gamma}_{k}=\frac{1}{JB}\sum_{i}^{JB}\gamma_{ik}$. By raising the forgetting factor to the power of $\hat{\gamma}_{k}$, components with higher expected responsibility receive more aggressive updates, whereas those with lower responsibilities retain a larger fraction of their previous statistics, thereby, preventing the sufficient statistics to be pushed towards zero.
Given the updated sufficient statistics, we proceed to the maximization step of the EM algorithm, wherein the GMM parameters $\Psi$ are updated. This step updates the mixture weights, means, and covariances by computing their maximum likelihood estimates based on the current sufficient statistics:
$\pi_{k}^{(t)}=\frac{S_{k,\pi}^{(t)}}{\sum_{k^{\prime}=1}^{K}S^{(t)}_{k^{\prime},\pi}},\quad\mu_{k}^{(t)}=\frac{S_{k,\mu}^{(t)}}{S_{k,\pi}^{(t)}},\quad\Sigma_{k}^{(t)}=\frac{S_{k,\Sigma}^{(t)}}{S_{k,\pi}^{(t)}}-\frac{S_{k,\mu}^{(t)}S_{k,\mu}^{(t)^{T}}}{S_{k,\pi}^{(t)}S_{k,\pi}^{(t)}}.$ (A.5)
In the context of our prototypical SSL framework, the updated means $\mu_{k}$ serve as prototypes for the latent assignments, i.e., $\mu_{k}=c_{k}$ for $k=1,\dots,K$. It is important to note that this maximization step is performed prior to the encoder’s optimizer step, ensuring that the prototypes reflect the most recent assignment statistics before any encoder parameters are updated.
We use a variation of split-and-merge (Ueda and Nakano, 2000) for regularization. But instead of merging low-weight clusters, we instead we apply a method we call split-resurrect. Whenever some component $k$ exceeds a weight threshold, we identify the lightest component $j$, reinitialize its mean (scale-aware random initialization) and reset its variance, and then split the mass of $k$ evenly, $\pi_{k}\leftarrow\pi_{j}\leftarrow\tfrac{1}{2}\pi_{k}^{\text{old}}$. For high cluster weights, we regularize the scale of the dominant mean by
$\hat{\mu}_{k}^{(t)}=\mu_{k}^{(t)}/\sqrt{\|\mu_{k}^{(t)}\|_{2}}$ (A.6)
which avoids dominant high-norm clusters in the GMM.
A.1 Ablation: Importance of the Different Components
We assess the impact of the individual components in our proposed decoupling strategy — the online Gaussian mixture. Specifically, we train CARP with a ViT backbone using 10 local crops on ImageNet-1k to evaluate each component. The linear accuracy is reported using PyTorch’s L-BFGS solver (Paszke et al., 2019; Liu and Nocedal, 1989).
As shown in Table A.1, the responsibility-forgetting mechanism introduced by Celaya and Agostini (2015) is essential given the sparse updates that occur when using a large number of prototypes. Without this mechanism, the forgetting factor $\eta$ drives the sufficient statistics toward zero, ultimately leading to collapse. While the remaining modifications are less detrimental, they nonetheless yield measurable gains in the encoder’s representational quality to varying degrees.
Table A.1: Ablation study of model components.
Responsibility Based Forgetting
(Celaya and Agostini, 2015)
Annealing
(Ueda and Nakano, 1998)
Resurrect
Rescaling
Lin.
0.1
71.65
71.97
71.99
72.12
72.25
Appendix B Training Details
B.1 Hyperparameters
For our experiments in Sections 4.3 and 4.4 our baselines DINO, DINO + KP and CARP are trained using a ViT-S/16 architecture while CARL is trained using a ResNet-50 backbone. Additionally, for CARL we increase the number of initialized prototypes to $65536$ to enable a fairer comparison and adopt the same training improvements incorporated in CARP, which includes multi-crop augmentation and added schedulers. Besides accommodating for these architectural changes in CARL and CARP, the rest of the official codebase is left untouched. For the DINO model we use the officially released codebase without applying any changes (Caron et al., 2021). For DINO+KP, we use the same codebase as for DINO with a minor modification: we incorporate the KoLeo-Prototype regularization proposed by Govindarajan et al. (2024), setting the regularization strength to the empirically validated value of $0.1$ reported in their paper.
We report the hyperparameters used to train the vision transformer backbones (Dosovitskiy et al., 2021) with CARP and CARP + Decoupling in Table B.1. In addition to these hyperparameters, CARP + Decoupling employs a weight threshold of $0.3$ for the resurrect strategy, which is relatively high given the $65536$ components in the mixture. Furthermore, for the forgetting factor $\eta$, we use a linear scheduler that increases from $0.1$ at the start of training to $0.5$ at its conclusion.
Table B.1: Training hyperparameters for ViT-Small configurations for CARP & CARP + Decoupling and ViT-Base configurations forCARP + Decoupling.
(a) ViT-Small Configuration
argument value
architecture vit_small
n_prototypes 65536
patch size 16
momentum_teacher 0.992
drop_path_rate 0.1
bottleneck_dim 256
global_crops_scale [0.32, 1.0]
local_crops_scale [0.05, 0.32]
optimizer AdamW
learning rate (lr) 55e4-4
weight_decay 0.04
weight_decay_end 0.4
freeze_prototypes false
min_lr 11e6-6
clip_grad 1.0
warmup_epochs 10
(b) ViT-Base Configuration
config value
architecture vit_base
n_prototypes 65536
patch size 16
momentum_teacher 0.996
drop_path_rate 0.1
bottleneck_dim 256
global_crops_scale [0.32, 1.0]
local_crops_scale [0.05, 0.32]
optimizer AdamW
learning rate (lr) 7.57.5e4-4
weight_decay 0.04
weight_decay_end 0.4
freeze_prototypes false
min_lr 22e6-6
clip_grad 0.3
warmup_epochs 10
B.1.1 Experiment: Uniqueness of prototypes under different thresholds
[…]
B.1.2 Experiment: Training Dynamics
[…]
B.1.3 Experiment: Effect of prototype diversity across head, medium and tail classes
We train all our baselines and CARP + Decoupling on iNaturalist2018 dataset using 10 local crops. The iNaturalist2018 challenge is a large scale species classification challenge that has long been the standard dataset for evaluating long-tailed performance. There are 8142 classes in this dataset, with 437,513 training images, and 24,426 validation images. We use the validation set for evaluation as the real test is deliberately left unavailable for public use. The dataset follows a heavy long-tail, with multiple classes appearing very few times in the training set. Following Liu et al. (2019), we consider all classes with more than 100 images each as head classes, between 20 to a 100 images each as medium classes, and less than 20 images each to be tail classes. Using this design choice, we have 842 head classes, 3701 medium classes, and 3599 tail classes.
We adopt the linear evaluation protocol of Caron et al. (2021) to assess all methods, with results reported in Table 2.
Appendix C Linear Evaluation
C.1 ImageNet-1k
[…]
Table C.1: ImageNet evaluation results for various prototypical SSL methods and architectures.
Method Backbone Dataset Epochs k-NN (%) Linear (%)
DINO (Caron et al., 2021) RN-50 ImageNet1kImageNet-1k 400 67.5 75.3
SWAV (Caron et al., 2020) RN-50 ImageNet1kImageNet-1k 400 65.0 74.6
CARP (Silva and Ramírez Rivera, 2023) RN-50 ImageNet1kImageNet-1k 400 67.7 75.3
CARP + Decoupling RN-50 ImageNet1kImageNet-1k 400 69.1 75.3
DINO (Caron et al., 2021) ViT-S/16 ImageNet1kImageNet-1k 300 72.8 76.2
iBOT (Zhou et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 300 74.6 77.4
CARP + Decoupling ViT-S/16 ImageNet1kImageNet-1k 300 74.1 76.2
MSN (Assran et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 600 76.9
iBOT (Zhou et al., 2022) ViT-S/16 ImageNet1kImageNet-1k 800 75.2 77.9
DINO (Caron et al., 2021) ViT-S/16 ImageNet1kImageNet-1k 800 74.5 77.0
DINO-vMF (Govindarajan et al., 2023) ViT-S/16 ImageNet1kImageNet-1k 800 74.7 77.0
iBOT-vMF (Govindarajan et al., 2024) ViT-S/16 ImageNet1kImageNet-1k 800 75.3 77.9
iBOT-vMF + KP (Govindarajan et al., 2024) ViT-S/16 ImageNet1kImageNet-1k 800 75.3 77.9
CARP + Decoupling ViT-S/16 ImageNet1kImageNet-1k 800 75.3 76.4
DINO (Caron et al., 2021) ViT-B/16 ImageNet1kImageNet-1k 400 76.1 78.2
DINO-vMF (Govindarajan et al., 2023) ViT-B/16 ImageNet1kImageNet-1k 400 77.4 78.8
iBOT (Zhou et al., 2022) ViT-B/16 ImageNet1kImageNet-1k 400 77.1 79.5
iBOT-vMF (Govindarajan et al., 2023) ViT-B/16 ImageNet1kImageNet-1k 400 78.7 80.3
iBOT-vMF + KP (Govindarajan et al., 2024) ViT-B/16 ImageNet1kImageNet-1k 400 78.8 80.5
CARP + Decoupling ViT-B/16 ImageNet1kImageNet-1k 400 76.7 78.1
C.2 iNaturalist2018
In Table C.2, we report the linear and fine-tuned classification accuracies on iNaturalist2018 for CARP + Decoupling trained with 10 local crops over 300 epochs. Baseline results are taken from Govindarajan et al. (2024).
Table C.2: iNat-2018 classification accuracies with full data. Results retrieved from Govindarajan et al. (2024).
Method Unique Protos. Linear (%) Fine-tuned (%)
ViT-Small/16
DINO-vMF (Govindarajan et al., 2023) 1380 $(2.1\%)$ 49.7 68.5
iBOT-vMF (Govindarajan et al., 2023) 1804 $(22.0\%)$ 50.1 69.4
iBOT-vMF (kd) (Govindarajan et al., 2023; Oquab et al., 2024) 1843 $(22.5\%)$ 50.5 69.1
iBOT-vMF (kp) (Govindarajan et al., 2024) 7895 $(96.4\%)$ 51.1 69.3
MSN ($\lambda=1$) (Assran et al., 2022) 3363 $(41.3\%)$ 53.8 63.5
PMSN ($\lambda=5$) (Assran et al., 2023a) 3005 $(36.9\%)$ 41.8 64.2
CARP + Decoupling 65536 $(100\%)$ 49.1 71.7
ViT-Base/16
iBOT-vMF (kd) (Govindarajan et al., 2023; Oquab et al., 2024) 1634 $(19.9\%)$ 50.4 73.3
iBOT-vMF (kp) (Govindarajan et al., 2024) 7573 $(92.4\%)$ 51.4 74.0
CARP + Decoupling 65536 $(100\%)$ 48.3 71.8